### v1

# “Quantum Geometric Nesting” and Solvable Model Flat-Band Systems

Abstract
We introduce the concept of “quantum geometric nesting” (QGN) to characterize the idealized ordering tendencies of certain flat-band systems implicit in the geometric structure of the flat-band subspace. Perfect QGN implies the existence of an infinite class of local interactions that can be explicitly constructed and give rise to solvable ground states with various forms of possible fermion bi-linear order, including flavor ferromagnetism, density waves, and superconductivity. For the ideal Hamiltonians constructed in this way, we show that certain aspects of the low-energy spectrum can also be exactly computed including, in the superconducting case, the phase stiffness. Examples of perfect QGN include flat bands with certain symmetries (e.g. chiral or time-reversal), and non-symmetry-related cases exemplified with an engineered model for pair-density-wave. Extending this approach, we obtain exact superconducting ground states with nontrivial pairing symmetry.
The properties of a correlated electronic system are largely determined by its electronic structure. One such familiar manifestation is the role of Fermi surface nesting conditions: if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e. $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order. If the overlap occurs under a momentum-space inversion about a certain wavevector $\bm{Q}/2$ (i.e. $\bm{k}\to\bm{Q}-\bm{k}$), the favored order is superconductivity (SC) with particle-particle (p-p) fermion bilinear order and a Cooper pair center of mass momentum $\bm{Q}$. (Such perfect nesting always exists at $\bm{Q}=0$ in time-reversal symmetric systems, where it leads to the familiar Cooper instability.) The thinking behind Fermi surface nesting is based on weak coupling mean-field analysis, in which ordering tendencies are diagnosed by divergent bare susceptibilities in corresponding channels. These concepts are indeed convenient and powerful (see e.g. Refs. Whangbo et al. (1991); Chuang et al. (2001); Kivelson et al. (2003); Shen et al. (2005); Lee et al. (2006); Johannes and Mazin (2008); Terashima et al. (2009); Ran et al. (2009); Nandkishore et al. (2012); Feng et al. (2013); Lee (2014); Jin et al. (2022) for typical applications). However, their validity is limited to weakly coupled systems, and they completely neglect another important piece of information in the electronic structure - the quantum geometry Resta (2011) and topology Bradlyn et al. (2017); Po et al. (2017); Kruthoff et al. (2017) encoded in the single-particle wavefunctions. An entirely different concept of “nesting” is thus needed for flat-band systems, whose bandwidths are small compared to the interaction strengths but which often have non-trivial quantum geometry and rich ordering phenomena Peotta and Törmä (2015); Tovmasyan et al. (2016); Kang and Vafek (2019); Rhim and Yang (2019); Hofmann et al. (2020); Bultinck et al. (2020); Kwan et al. (2021a); Khalaf et al. (2021); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Kwan et al. (2021b); Vafek and Kang (2021); Peri et al. (2021); Liu et al. (2021); Rossi (2021); Angeli and MacDonald (2021); Wang et al. (2022); Hu et al. (2022); Huhtinen et al. (2022); Törmä et al. (2022); Chan et al. (2022a); Herzog-Arbeitman et al. (2022a); Mera and Mitscherling (2022); Pan et al. (2022); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Lin et al. (2022); Herzog-Arbeitman et al. (2022a); Peotta (2022); Khalaf et al. (2022); Chan et al. (2022b); Kang and Vafek (2019); Hofmann et al. (2023); Wan et al. (2023); Swaminathan et al. (2023); Hu et al. (2023); Löthman and Black-Schaffer (2017); Chen and Law (2023); Iskin (2023a); Peotta et al. (2023); Iskin (2023b); Tam and Peotta (2023); Zhang et al. (2023); Pan et al. (2023); Crépel et al. (2023); Mao et al. (2023); Kwan et al. (2023); Sahay et al. (2023); Jiang and Barlas (2023); Scheer and Lian (2023a, b); Devakul et al. (2023); Herzog-Arbeitman et al. (2023); Jia et al. (2023); Wu et al. (2023); Chen and Huang (2023).
In this paper, we propose such a concept, which we call “Quantum Geometric Nesting” (QGN), that focuses on the geometric aspects of the electronic structure in flat-band systems. In contrast to Fermi surface nesting, here we start from a strong-coupling perspective in which the band dispersion is assumed negligible in comparison with the interaction strength, and then ask: for a given quantum geometry of the flat bands, are there any natural forms of order? If so, what sort of interactions favor them? We will show that these questions can be addressed by examining whether a set of QGN conditions is satisfied. For “perfectly nested quantum geometry,” we construct ideal Hamiltonians containing arbitrary combinations of an infinite class of engineered short-range two-body interactions that have solvable ground states with fermion bilinear order of a character determined by the quantum geometry. For these ideal Hamiltonians, the few-body excitations are solvable as well, including single-particle (quasi-particle), particle-hole (exciton), and two-particle (Cooper pair) excitations. For the p-p case, we further show that these ideal Hamiltonians host a (fine-tuned) pseudospin $SU(2)$ symmetry, whose stiffness can be inferred from the excitation spectra. These solvable interacting Hamiltonians can either directly describe realistic systems or serve as starting points for various perturbative studies of additional terms. This type of intrinsically strong coupling viewpoint has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al. (2016); Kwan et al. (2021b); Vafek and Kang (2021); Kang and Vafek (2019); Khalaf et al. (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021); Pan et al. (2022); Lin et al. (2022); Herzog-Arbeitman et al. (2022a); Zhang et al. (2022); Herzog-Arbeitman et al. (2022b); Crépel et al. (2023); Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023); Kwan et al. (2023), and it is remarkable that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems. More possibilities also exist within the framework of QGN, which are exemplified with an engineered model for pair-density-wave. Extending the construction scheme, we further obtain an SSH-type model with solvable $p$-wave pairing groundstates.
We will present the conclusions in the main text and defer the detailed proofs and technical discussions to the corresponding sections in the Supplemental Materials (SM). In Fig. 1, we provide a schematic flowchart for the central results as a guide for readers.
Figure 1: Schematic flowchart illustrating the structure of the theory: Given any electronic structure containing isolated flat bands, one can test whether a QGN condition is satisfied; if it is, one or more fermion bilinear order parameters (OP) will be suggested (Sec. I). For each OP, infinitely many “ideal” interacting Hamiltonians can be systematically constructed (Sec. II), which, after projecting onto the flat-band Hilbert space $\mathcal{H}_{\text{flat}}$, feature solvable ground states with the corresponding order (Sec. III) and solvable aspects of the collective excitation spectrum (Sec. IV). Some examples of electronic structures with perfect QGN can be found in Sec. V. Possible extensions of this approach are commented on in Sec. VI.
I Formalism
I.1 Preliminaries
We consider an electronic structure encoded in a tight-binding model with charge conservation and translation symmetry among $\mathsf{V}$ unit cells in arbitrary dimension:
$\hat{H}_{0}=\sum_{\bm{R}\bm{R}^{\prime},\mu\nu}t_{\mu\nu}(\bm{R}-\bm{R}^{\prime})\ \hat{c}^{\dagger}_{\bm{R},\mu}\hat{c}_{\bm{R}^{\prime},\nu}$ (1)
where $\bm{R}$ indicates the unit cell position, Greek letters in lowercase $\mu,\nu,\dots$ denote a “flavor” index which, for simplicity, includes both orbital and spin indices unless otherwise specified. The annihilation operators for Bloch states with a given flavor index are expressed as $\hat{c}_{\bm{k},\mu}\equiv\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{R}}\hat{c}_{\bm{R},\mu}/\sqrt{\mathsf{V}}$ (periodic gauge) while the single particle Hamiltonian is diagonalized in a band basis,
$\hat{H}_{0}$ $=\sum_{\bm{k},n}\epsilon_{n}(\bm{k})\hat{\gamma}^{\dagger}_{\bm{k},n}\hat{\gamma}_{\bm{k},n},\quad\hat{\gamma}_{\bm{k},n}=U_{n\mu}^{\dagger}(\bm{k})\hat{c}_{\bm{k},\mu},$ (2)
where $U_{n\mu}^{\dagger}(\bm{k})$ are unitary matrices and we use Latin letters in lowercase ($n,m,\dots$) to label the bands. We assume there are $\mathsf{N}$ flavors and thus $\mathsf{N}$ bands in total.
We will quantify the degree of QGN for flat-band systems by utilizing the geometric information in $U_{\mu n}(\bm{k})$. To do so, we consider an idealized setup where there are $\mathsf{N}_{\text{flat}}$ nearly degenerate and flat bands isolated from all the other bands by a large gap $\Delta$, such that the relevant interaction strength $V$ is small compared to $\Delta$ but large compared to the flat bands’ bare bandwidth $W$,
$W\ll V\ll\Delta.$ (3)
Without loss of generality, we can relabel all the bands to make the $n=1\dots\mathsf{N}_{\text{flat}}$ bands the flat bands. We then define the $\bm{k}$-dependent projection matrix for the flat bands
$P_{\mu\nu}(\bm{k})\equiv\sum_{n\leq\mathsf{N}_{\text{flat}}}U_{\mu n}(\bm{k})U_{n\nu}^{\dagger}(\bm{k})$ (4)
and, for later convenience, the projection matrix onto the remaining bands, $Q_{\mu\nu}(\bm{k})\equiv\delta_{\mu\nu}-P_{\mu\nu}(\bm{k})$.
For the ideal scenario we are assuming, we define the “flat-band subspace”, $\mathcal{H}_{\text{flat}}$, as the tensor product of the Fock space of the flat bands and a “vacuum” state of the other modes $\gamma_{\bm{k},n=\mathsf{N}_{\text{flat}}+1\dots\mathsf{N}}$ (i.e. all the modes with energy higher or lower than the flat bands are respectively empty or occupied). We formally define the projector onto $\mathcal{H}_{\text{flat}}$ as $\hat{P}$. To leading order of the perturbation series treating both $V/\Delta$ and $W/V$ as small parameters, we can set $\hat{H}_{0}=0$ and project the interactions, $\hat{H}_{\text{int}}$, onto $\mathcal{H}_{\text{flat}}$. Formally, this means studying the effective Hamiltonian $\hat{P}\hat{H}_{\text{int}}\hat{P}$, but operationally this can be more conveniently done by retaining terms that only involve $\hat{\gamma}_{\bm{k},n\leq\mathsf{N}_{\text{flat}}}$, or equivalently, by replacing all $\hat{c}_{\bm{k}\mu}\rightarrow P_{\mu\nu}(\bm{k})\hat{c}_{\bm{k}\nu}$ in the interacting Hamiltonian  [^1] We note that this scheme in general will result in some additional quadratic terms when compared to $\hat{P}\hat{H}_{\text{int}}\hat{P}$, which can be operationally viewed as part of $\hat{H}_{0}$. See Sec. II below and Sec. II D of SM for detailed discussions..
I.2 General Definitions
In terms of the gauge-invariant matrices $P$ and $Q$, the key information of QGN at a given wavevector $\bm{Q}$ for p-p or p-h channel is encoded in the linear operator
$\Pi^{\text{p-p},\bm{Q}}_{\mu^{\prime}\nu^{\prime};\mu\nu}=\frac{1}{\mathsf{V}}\sum_{\bm{k}}\left[P^{\star}_{\mu^{\prime}\mu}(\frac{\bm{Q}}{2}+\bm{k})Q_{\nu\nu^{\prime}}(\frac{\bm{Q}}{2}-\bm{k})+(P\leftrightarrow Q)\right],$ (5)
$\Pi^{\text{p-h},\bm{Q}}_{\mu^{\prime}\nu^{\prime};\mu\nu}=\frac{1}{\mathsf{V}}\sum_{\bm{k}}\left[P_{\mu^{\prime}\mu}(\bm{k}+\frac{\bm{Q}}{2})Q_{\nu\nu^{\prime}}(\bm{k}-\frac{\bm{Q}}{2})+(P\leftrightarrow Q)\right].$ (6)
Viewed as a matrix (with $\mu\nu$ a grouped index and $\mu^{\prime}\nu^{\prime}$ the other), $\Pi$ is hermitian and positive semi-definite and can be diagonalized with non-negative eigenvalues.
By definition, an electronic structure has “perfect QGN” at a wavevector $\bm{Q}$ in the p-p or p-h channel when the corresponding $\Pi$ has eigenvalue(s) equal to 00. In other words, a perfect QGN is satisfied if there exist a $\mathsf{N}\times\mathsf{N}$ matrix $N^{\bm{Q}}_{\mu\nu}$, which we call the “nesting matrix”, satisfying (for any $\mu^{\prime}\nu^{\prime}$)
$\sum_{\mu\nu}\Pi_{\mu^{\prime}\nu^{\prime};\mu\nu}^{\bm{Q}}N^{\bm{Q}}_{\mu\nu}=0.$ (7)
We will show that the favored OP operator within the flat-band subspace can be expressed in terms of $N^{\bm{Q}}$ as:
$\hat{O}^{\text{p-p},\bm{Q}}$ $\equiv\frac{1}{\mathsf{V}}\sum_{\bm{k};n,m\leq\mathsf{N}_{\text{flat}}}F^{\text{p-p},\bm{Q}}_{nm}(\bm{k})\hat{\gamma}_{\bm{Q}/2+\bm{k},n}\hat{\gamma}_{\bm{Q}/2-\bm{k},m}$ (8)
$\hat{O}^{\text{p-h},\bm{Q}}$ $\equiv\frac{1}{\mathsf{V}}\sum_{\bm{k};n,m\leq\mathsf{N}_{\text{flat}}}F^{\text{p-h},\bm{Q}}_{nm}(\bm{k})\hat{\gamma}^{\dagger}_{\bm{k}+\bm{Q}/2,n}\hat{\gamma}_{\bm{k}-\bm{Q}/2,m}$ (9)
where $F$ is the form factor
$F^{\text{p-p},\bm{Q}}_{nm}(\bm{k})\equiv\sum_{\mu\nu}U_{\mu n}(\frac{\bm{Q}}{2}+\bm{k})N^{\text{p-p},\bm{Q}}_{\mu\nu}U_{\nu m}(\frac{\bm{Q}}{2}-\bm{k})$ (10)
$F^{\text{p-h},\bm{Q}}_{nm}(\bm{k})\equiv\sum_{\mu\nu}U^{\dagger}_{n\mu}(\bm{k}+\frac{\bm{Q}}{2})N^{\text{p-h},\bm{Q}}_{\mu\nu}U_{\nu m}(\bm{k}-\frac{\bm{Q}}{2}).$ (11)
One important property (which can serve as an equivalent definition) of QGN is that for arbitrary $\bm{k}$, the form factor $F_{nm}(\bm{k})$ defined by a nesting matrix is a block-diagonal matrix that disconnects flat band indices from the others, i.e. $F_{nm}(\bm{k})=0$ as long as $n,m$ are not simultaneously $\leq\mathsf{N}_{\text{flat}}$ or $>\mathsf{N}_{\text{flat}}$.
For QGN to be non-trivial in the p-p channel, it is additionally required that $N^{\bm{Q}}$ cannot be symmetric since $\hat{O}^{\text{p-p},\bm{Q}}$ vanishes in this case. (It is proven by explicit construction that valid, anti-symmetric solutions can indeed exist for $\Pi$.) On the other hand, the identity matrix is always an allowed solution for Eq. 7 in the p-h case for $\bm{Q}=\bm{0}$, but this is a trivial one in the sense that the resulting $\hat{O}^{\text{p-h},\bm{Q}=\bm{0}}$ is simply the density operator not associated with any broken symmetry. Excluding all the unwanted solutions, for a generic flat-band system, the value of the smallest eigenvalue of $\Pi^{\bm{Q}}$, $\varpi_{0}^{\bm{Q}}$, is a natural dimensionless measure of “nestability” that quantifies the closeness of a system to perfect QGN in the corresponding channel at $\bm{Q}$  [^2] We note that these concepts are spiritually similar to the “ideal band condition” Roy (2014); Claassen et al. (2015); Mera and Ozawa (2021); Ledwith et al. (2022); Wang et al. (2021); Wang and Liu (2022); Wang et al. (2023); Dong et al. (2023) and “vortexability” Ledwith et al. (2023) which are concerned with whether a Chern band is suitable for hosting fractional quantum Hall states.. Later in Sec. V we will give several examples of electronic structures with perfect QGN, i.e. $\varpi_{0}^{\bm{Q}}=0$.
Depending on the channel, momentum $\bm{Q}$, and nesting matrix $N^{\bm{Q}}$, Eq. 8 or 9 represent fermion bilinear orders with different natures. In the p-p channel, if $\bm{Q}=\bm{0}$, the order is a uniform SC order; while for $\bm{Q}\neq\bm{0}$ it is pair-density-wave (PDW, not to be confused with the other DWs in the p-h channel) order of the single-$\bm{Q}$, Fulde–Ferrell type Agterberg et al. (2020). In the p-h channel, if $\bm{Q}=\bm{0}$, non-trivial QGN ($N^{\text{p-h},\bm{0}}\neq\mathbbm{1}$) leads to flavor-symmetry-breaking orders, including flavor ferromagnetism and inter-flavor coherent states; the $\bm{Q}\neq\bm{0}$ case promotes those orders to various forms of flavor DWs (e.g. spin DWs), with an additional possibility of pure charge DW order for $N^{\text{p-h},\bm{Q}}=\mathbbm{1}$.
A system can satisfy multiple QGN conditions in different channels and/or at different $\bm{Q}$s. It is also possible that there can be more than one linearly independent nesting matrix for each channel and $\bm{Q}$, which typically signals a further internal symmetry breaking, e.g. triplet SC in spin invariant systems.
II Ideal interacting Hamiltonians
We now present a general strategy to construct a solvable “ideal” interacting Hamiltonian, $\hat{H}_{\text{int}}$, for a flat-band system with perfect QGN. It is an engineered model in the sense that it targets the OP $\hat{O}$ defined as in Eq. 8 or 9; specifically, it is constructed to have ground states that exhibit a pattern of spontaneous symmetry breaking associated with an anomalous expectation value of $\hat{O}$. For reasons that will be made clear in the next section, we write $\hat{H}_{\text{int}}$ in the form:
$\hat{H}_{\text{int}}=\sum_{\bm{R}\bm{R}^{\prime};IJ}V_{IJ}(\bm{R}-\bm{R}^{\prime})\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}_{\bm{R}}^{(I)}\rangle\right)\left(\hat{S}^{(J)}_{\bm{R}^{\prime}}-\langle\hat{S}_{\bm{R}^{\prime}}^{(J)}\rangle\right)$ (12)
where $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ is any function whose Fourier transform
$V_{IJ}(\bm{q})\equiv\frac{1}{\mathcal{V}}\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{q}\cdot\bm{R}}\ V_{IJ}(\bm{R})$ (13)
is positive semi-definite for all $\bm{q}$. Thus, by construction, the eigen-energies of $\hat{H}_{\text{int}}$ are non-negative. $\{\hat{S}^{(I)}_{\bm{R}}\}$ is an infinite set of linearly independent Hermitian operators centered in any unit cell $\bm{R}$, which, after projection onto $\mathcal{H}_{\text{flat}}$, commute with $\hat{O}$ (in Eq. 8/9):
$\hat{P}\ [\hat{S}^{(I)}_{\bm{R}},\hat{O}]\ \hat{P}=0\ .$ (14)
These hermitian operators, $\hat{S}^{(I)}_{\bm{R}}$, are labeled with Latin letters in uppercase, and $\langle\hat{S}_{\bm{R}}^{(I)}\rangle$ are the corresponding ground-state expectation values. It should be noted that these operators are defined in the entire Hilbert space, whereas the above commutation relation only holds after projection.
For generic flat-band systems, no set of local hermitian operators that satisfy Eq. 14 exists. However, when $N$ and $U$ in Eq. 8/9 derive from a band geometry with perfect QGN, there is an infinite class of solutions that take a special flavor-momentum separable form
$\hat{S}^{(I)}_{\bm{R}}=\sum_{\bm{R}_{1}\bm{R}_{2};\mu\nu}s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}B^{(I)}_{\mu\nu}\hat{c}^{\dagger}_{\bm{R}+\bm{R}_{1},\mu}\hat{c}_{\bm{R}+\bm{R}_{2},\nu},$ (15)
where $B^{(I)}$ is a hermitian matrix and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=[s^{(I)}_{\bm{R}_{2},\bm{R}_{1}}]^{\star}$ is the spatial coeffient. In order for Eq. 14 to hold, the constraints on the coefficient $s^{(I)}$ and matrix $B^{(I)}$ are rather loose, in that they only need to obey
$s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=\left[s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}\right]^{\star}\mathrm{e}^{-\mathrm{i}(\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}},$ (16)
$N^{\text{p-p},\bm{Q}}\cdot B^{(I)}+[B^{(I)}]^{T}\cdot N^{\text{p-p},\bm{Q}}=0,$ (17)
for the p-p case, and
$s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=0\ \ \text{if}\ \ (\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}\neq 0\mod 2\pi,$ (18)
$N^{\text{p-h},\bm{Q}}\cdot B^{(I)}-B^{(I)}\cdot N^{\text{p-h},\bm{Q}}=0.$ (19)
for the p-h case. We derive in Sec. II B&C of SM a systematic way of constructing all possible $B^{(I)}$ and prove that there are at least $3\mathsf{N}_{\text{flat}}$ of them in the p-p case, and $\mathsf{N}_{\text{flat}}$ of them in the p-h case when $2\bm{Q}=\bm{0}$. (Some explicit worked examples are presented in Sec. V.) In particular, the identity matrix is always a solution for $B$ in the p-h case, which makes $\hat{S}_{\bm{R}}$ a local density operator for the unit cell at $\bm{R}$. Clearly, even given all the constraints, we are still left with an infinite class of possible interaction terms with considerable leeway, at least in choosing $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}$, which thus could substantially overlap with realistic interactions.
One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”. For the special case of Hubbard interactions, a property called the “uniform pairing condition” Tovmasyan et al. (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al. (2022a). It is possible to generalize this condition, which we will exemplify in Sec. V.3. However, these quadratic terms are not of central focus since the current work is concerned with the existence of ideal interacting Hamiltonians for a given set of isolated flat bands, whose quantum geometry is encoded in the projection matrix and is unaffected by these terms to the leading order in $V/\Delta$. An operational workaround is to simply regard these additional quadratic terms as parts of $\hat{H}_{0}$. For more detailed discussions about this nomenclature subtlety, see Sec. II D of SM.
III Solvable ground state
In this section, we present a general recipe for obtaining the many-body ground states for an ideal Hamiltonian $\hat{H}_{\text{int}}$ constructed in the previous section. We first define a “pseudo-Hamiltonian” within $\mathcal{H}_{\text{flat}}$ with $\theta$ a free $U(1)$ phase
^$\hat{\mathcal{E}}=\mathsf{V}(\mathrm{e}^{\mathrm{i}\theta}\hat{O}+\mathrm{e}^{-\mathrm{i}\theta}\hat{O}^{\dagger}),$ (20)
which can be viewed as a trial Hamiltonian of the original system. Since $\hat{\mathcal{E}}$ commutes with the ideal $\{\hat{S}^{(I)}_{\bm{R}}\}$ and thus the ideal $\hat{H}_{\text{int}}$ within $\mathcal{H}_{\text{flat}}$, any non-degenerate eigenstate of $\hat{\mathcal{E}}$, $|\Psi\rangle$, must also be an eigenstate of $\{\hat{S}^{(I)}_{\bm{R}}\}$ and $\hat{H}_{\text{int}}$ projected onto $\mathcal{H}_{\text{flat}}$. Taking $\langle\hat{S}_{\bm{R}}^{(I)}\rangle\equiv\langle\Psi|\hat{S}^{(I)}_{\bm{R}}|\Psi\rangle$ in Eq. 12 implies $|\Psi\rangle$ is a zero-energy eigenstate of $\hat{H}_{\text{int}}$ within $\mathcal{H}_{\text{flat}}$ and thus a ground state.
Therefore, the task is to find the unique eigenstates of $\hat{\mathcal{E}}$ and show that they are ordered. Since $\hat{\mathcal{E}}$ is quadratic in fermion operators, its spectrum and eigenmodes can be easily solved. In general, the bands of $\hat{\mathcal{E}}$ are flat, and the values of the “pseudo-energies” are determined by $N^{\bm{Q}}$.
If the QGN is in the p-p channel, the spectrum of $\hat{\mathcal{E}}$ is given by $\mathsf{N}_{\text{flat}}$ Bogoliubov-de-Gennes quasiparticle pseudo-bands with perfect particle-hole symmetry (since no $\hat{\gamma}^{\dagger}\hat{\gamma}$ term is present). Fully occupying all the negative pseudo-energy bands gives rise to a state $|\Psi_{\text{SC}}(\theta)\rangle$ with eigenvalue $\Phi_{0}$, which is a sum of all the pseudo-energies of the occupied bands and thus is non-zero and extensive. $\Phi_{0}$ is a non-degenerate eigenvalue of $\hat{\mathcal{E}}$, since any change of the occupation configuration of the BdG quasi-particles will also change the eigenvalue  [^3] For the clarity of key points, here we focus on the case where $N$ is full-rank. The discussions on the general cases can be found in Sec. III of SM.. Thus, $|\Psi_{\text{SC}}(\theta)\rangle$ is a ground-state of $\hat{H}_{\text{int}}$, and it hosts off-diagonal-long-range-order (ODLRO) Yang (1962) with a superconducting phase $\theta$ in the sense that
$\langle\Psi_{\text{SC}}(\theta)|\hat{O}|\Psi_{\text{SC}}(\theta)\rangle=\frac{\Phi_{0}}{2\mathsf{V}}\mathrm{e}^{-\mathrm{i}\theta}\equiv\frac{\phi_{0}}{2}\mathrm{e}^{-\mathrm{i}\theta}$ (21)
Note that by doing a global gauge transformation $\hat{\gamma}\rightarrow\hat{\gamma}\mathrm{e}^{-\mathrm{i}\theta/2}$, we are able to transform ^$\hat{\mathcal{E}}(\theta)\rightarrow\hat{\mathcal{E}}(\theta=0)$ and thus $|\Psi_{\text{SC}}(\theta)\rangle\rightarrow|\Psi_{\text{SC}}(\theta=0)\rangle$, while keeping $\hat{S}^{(I)}_{\bm{R}}$ unchanged. Thus, all $|\Psi_{\text{SC}}(\theta)\rangle$ with different $\theta$ are simultaneous eigenstate of $\hat{S}^{(I)}_{\bm{R}}$ with the same eigenvalue. The corresponding ground state of $\hat{H}_{\text{int}}$ with a fixed (even) particle number $\mathsf{N}_{e}$ can be constructed as
$|\Psi_{\text{SC}}(\mathsf{N}_{e})\rangle\equiv\int_{0}^{2\pi}\mathrm{e}^{-\mathrm{i}\theta\mathsf{N}_{e}/2}|\Psi_{\text{SC}}(\theta)\rangle\ .$ (22)
In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector. It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure. Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice. In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
While the four-fermion term in $H_{\text{int}}$ is translationally invariant, the single particle term is only translationally invariant when all $\langle\hat{S}^{(I)}_{\bm{R}}\rangle$ are. For any order with $\bm{Q}=\bm{0}$, the translational invariance of $\langle\hat{S}^{(I)}_{\bm{R}}\rangle$ is obvious. Even for $\bm{Q}\neq\bm{0}$, $\langle\hat{S}^{(I)}_{\bm{R}}\rangle$ can still be translationally invariant. For example, because the p-p orders we have constructed are of the FF type, translation by $\bm{R}$ is equivalent to a phase change of the OP by $\theta\to\theta+\bm{Q}\cdot\delta\bm{R}$; since $\hat{S}^{(I)}_{\bm{R}}$, by assumption, are gauge invariant, their expectation values remain constant under translation. Similar situations can arise in the p-h cases, for example, when the order involved is a generalized “spiral order” and the used $\hat{S}^{(I)}_{\bm{R}}$ only overlap with the magnitude of the OP. In general, however, certain choices of $\hat{H}_{\text{int}}$ may explicitly break translational symmetry, but still preserve other symmetries that are broken spontaneously by $|\Psi\rangle$.
IV Excitations in the ideal models
In the ideal models constructed in Sec. II, aspects of the excitation spectrum can also be determined in a systematic fashion, which we discuss in this section.
IV.1 Few-body excitations
The strategy to solve the excitations with simple compositions is to construct operators $\hat{\xi}^{(c)}_{\bm{P}}$ within each charge-$c$, momentum-$\bm{P}$ sector that satisfy:
$[\hat{H}_{\text{int}},\hat{\xi}^{(c)}_{\bm{P}}]|\Psi\rangle=E^{(c)}(\bm{P})\hat{\xi}^{(c)}_{\bm{P}}|\Psi\rangle.$ (23)
This is easier said than done for typical interacting Hamiltonians, where the commutator of a one-particle operator, $\gamma_{a}^{\dagger}$, with the Hamiltonian produces higher order terms such as $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$, and the commutator of many-particle operators produces still higher order terms in a hierarchy that never closes. However, because $|\Psi\rangle$ is a zero-mode of all $\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}^{(I)}_{\bm{R}}\rangle\right)$ within $\mathcal{H}_{\text{flat}}$, sectors with e.g. $\gamma_{a}^{\dagger}$ and $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$ excitations cannot mix Kang and Vafek (2019); Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a); Vafek and Kang (2021) and thus the composition does not changed after the commutation. For example, for two-particle (Cooper pair) excitations,
$[\hat{H}_{\text{int}},\hat{\gamma}^{\dagger}_{\bm{k}^{+},n}\hat{\gamma}^{\dagger}_{\bm{k}^{-},m}]|\Psi\rangle$
$=$ $\sum_{\bm{k}^{\prime},n^{\prime}m^{\prime}\leq\mathsf{N}_{\text{flat}}}\Gamma^{\text{p-p},\bm{P}}_{nm,\bm{k};n^{\prime}m^{\prime},\bm{k}^{\prime}}\hat{\gamma}^{\dagger}_{\bm{k}^{\prime+},n^{\prime}}\hat{\gamma}^{\dagger}_{\bm{k}^{\prime-},m^{\prime}}|\Psi\rangle$ (24)
where $\bm{k}^{\pm}\equiv\bm{P}/2\pm\bm{k}$. The commutation relations for single-particle/hole, two-hole, and particle-hole (exciton) excitations can be similarly derived (see Sec. IV A-C of SM for computationally feasible expressions in terms of parameters in the Hamiltonian). By diagonalizing the scattering matrix $\Gamma$, one can compute the corresponding eigenstates and spectrum of an excitation Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a). For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian. These excitations can be computed with polynomial scaling in the system size, rather than exponentially as in most interacting systems, and in this sense are solvable. For certain cases when the form of interaction is simple (e.g. attractive Hubbard Herzog-Arbeitman et al. (2022a)), the quasi-particle spectrum can be even analytically solvable.
IV.2 Pseudo-spin $SU(2)$ symmetry of the p-p case and the phase stiffness
In general, different types of excitations do not have any relation. However, in any ideal model in the p-p case, some of the excitations with the same fermion parity but different charge can be related, since $\hat{O}^{\text{p-p},\bm{Q},\dagger}$ or $\hat{O}^{\text{p-p},\bm{Q}}$ adds charge ±$\pm 2$ and momentum ±$\pm\bm{Q}$ with zero energy cost,
$\hat{\xi}^{(c\pm 2)}_{\bm{P}\pm\bm{Q}}$ $\equiv[\hat{\xi}^{(c)}_{\bm{P}},\hat{O}^{\text{p-p},\bm{Q},\dagger}\text{ or }\hat{O}^{\text{p-p},\bm{Q}}]$ (25)
creates a charge $c\pm 2$, momentum $\bm{P}\pm\bm{Q}$, energy $E^{(c\pm 2)}(\bm{P}\pm\bm{Q})=E^{(c)}(\bm{P})$ excitation when acting on a ground state. Therefore, for the p-p case, the particle and hole excitations share the same spectrum, and certain particle-particle, particle-hole, and hole-hole pair excitations share the same spectra. Furthermore, lowest branch of $E^{(0,\pm 2)}(\bm{P})$ must hit zero at $\bm{P}=0,\pm\bm{Q}$ since $\hat{O}^{\text{p-p},\bm{Q},\dagger}$ creates a zero-energy p-p excitation with momentum $\bm{Q}$. This is due to a pseudospin $SU(2)$ symmetry in the p-p case, which is generated by:
$\hat{J}^{x,y}$ $\equiv(\hat{O}^{\text{p-p},\dagger}\pm\hat{O}^{\text{p-p}})/2,\quad\hat{J}^{z}\equiv[\hat{O}^{\text{p-p},\dagger},\hat{O}^{\text{p-p}}]/4$ (26)
where for clarity, we consider the single-component SC case; multicomponent generalizations are discussed in Sec. IV D of SM. Clearly, the SC ground state $|\Psi_{\text{SC}}\rangle$ is a form of pseudospin ferromagnetism. Thus, by comparing the excitation spectrum to that of a non-linear sigma model of broken $SU(2)$ symmetry Watanabe and Murayama (2014), we obtain the following estimate for the pseudospin stiffness tensor:
$\kappa_{ij}=\phi_{0}M^{-1}_{ij}$ (27)
where $M^{-1}_{ij}$ is the Cooper pair inverse mass tensor, extracted by a Taylor expansion of the lowest branch of the charge-$2$ spectrum $E^{(2)}(\bm{P}=\bm{Q}+\delta\bm{Q})\approx\frac{1}{2}\delta\bm{Q}_{i}M^{-1}_{ij}\delta\bm{Q}_{j}$, and $\phi_{0}$ is the ODLRO density defined in Eq. 21.
Given the quadratic dispersion of the Nambu-Goldstone mode, the Mermin-Wagner-Coleman theorem implies that spontaneous breaking of the $SU(2)$ symmetry cannot occur at finite temperature when the space dimension is $d\leq 2$. (At zero temperature, in contrast, ferromagnetism is allowed in low dimensions.) However, we note that the susceptibility $\chi(T)\sim T^{3}/|\kappa|^{4}\exp(4\pi|\kappa|/T)$ is large at low temperature $T\rightarrow 0$ Zinn-Justin (2021), where $|\kappa|\equiv\sqrt[d]{\det{\kappa}}$ is the geometric mean of the stiffness in different directions. Then, when $d=2$, certain explicit breaking of the symmetry down to charge $U(1)$ will result in a quasi-long-range ordering of the $U(1)$ phase with a Berezinskii–Kosterlitz–Thouless (BKT) transition temperature $T_{c}$ that can be estimated as
$\delta\cdot\chi(T_{c})\sim 1\implies T_{c}\sim 4\pi|\kappa|/\ln(|\kappa|/\delta)$ (28)
which quickly becomes comparable to $|\kappa|$ for non-zero symmetry-lowering energy scale $\delta$. We note that multiple factors can lead to a deviation from the idealized setup we constructed, and thus a weak explicit lowering of the pseudospin $SU(2)$ symmetry, including (but not limited to) the finite band gap, imperfect QGN and flatness of the bands, and the deviation of the interactions from the ideal form. Depending on the nature of those perturbations, the superconducting state may remain phase-coherent up to a temperature comparable to $|\kappa|$ for all dimensions $d\geq 2$. Our approach provides a direct estimate of the superfluid stiffness of nearly ideal systems, providing complementary estimates to generic upper bounds Hazra et al. (2019); Hofmann et al. (2022); Mao and Chowdhury (2023a, b).
V Examples of electronic structures with perfect QGN
V.1 Spinful time-reversal symmetric flat bands
We first note that QGN can be guaranteed by symmetries. The simplest example of this is the combination of spin $S^{z}$ conservation and time-reversal symmetry. In such a system, we recognize the Wannier degeneracy (separating out the spin index $s$ from the orbital index $\mu$):
$U_{\mu s,ns^{\prime}}(\bm{k})$ $=\delta_{ss^{\prime}}U_{\mu n}^{s}(\bm{k})\ ,\ \ U_{\mu n}^{s}(\bm{k})=\left[U_{\mu n}^{\bar{s}}(-\bm{k})\right]^{\star}$ (29)
Then one can verify $N^{\text{p-p},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}\equiv\sigma^{y}_{ss^{\prime}}\mathbbm{1}_{\mu\nu}$ satisfies the perfect QGN condition in the p-p channel with $\bm{Q}=\bm{0}$. Solvable Hamiltonians can thus be constructed with $B^{(I)}_{\mu s,\nu s^{\prime}}$ matrices in Eq. 15 taking the form $\sigma^{\mathsf{i}=0,x,y,z}_{ss^{\prime}}\tilde{B}^{(I)}_{\mu\nu}$, where $\sigma^{\mathsf{i}}$ are spin Pauli matrices and $\tilde{B}^{(I)}$ are arbitrary symmetric (for $\mathsf{i}=x,y,z$) or antisymmetric (for $\mathsf{i}=0$) hermitian matrices for the orbital indices. This includes a huge class of possible interactions, e.g. the attractive Hubbard interactions on each site and orbital, whose solvability was pointed out in Refs. Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al. (2023); Liang et al. (2017); Verma et al. (2021); Hazra et al. (2018); Xie et al. (2020); Herzog-Arbeitman et al. (2022b); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022a).
V.2 Chiral symmetric flat bands
A less obvious example is associated with chiral symmetry. Specifically, when the lattice is bipartite and there are degenerate flat bands, one can verify that $N^{\text{p-h},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}=\delta_{\mu\nu}(-1)^{\mu}\tilde{N}_{ss^{\prime}}$ is a nesting matrix for the p-h channel at $\bm{Q}=\bm{0}$, where the parity of the orbital index $\mu$ represents the sublattice group, and $\tilde{N}$ is an arbitrary hermitian matrix associated with the conserved flavor index $s$ (e.g. spin and/or valley). The condensation of the corresponding OP then leads to symmetry breaking of the conserved flavor symmetry with a pattern specified by $\tilde{N}$ (e.g. spin/valley polarization, inter-valley coherence, …). This is exactly the case extensively discussed in the studies on magic-angle twisted bilayer graphene in its chiral-flat limit Kang and Vafek (2019); Vafek and Kang (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021). We also note that there is a large class of electronic structures that feature flat bands protected by chiral symmetry Călugăru et al. (2022) and are approximately realized in materials Regnault et al. (2022).
It is of particular interest to identify and predict what interactions could lead to certain specific symmetry-breaking patterns in flat band materials. In this regard, the construction of solvable models for different OPs can provide important information by suggesting interactions that only favor certain types of symmetry-breaking patterns. We leave material-specific investigations to future studies.
V.3 Non-symmetry-related cases: An engineered model for pair-density-wave
It is probably even more interesting to study cases in which perfect QGN is not associated with symmetry. To show such a possibility, we construct a simple model of a dimerized bilayer on a square lattice in which $\tau^{0,x,y,z}$ are Pauli matrices acting on orbital/layer indices, and $s$ is the spin index,
$\hat{H}_{0}=\sum_{\bm{k},s,\mu\nu}\left[2t\left(\cos k_{x}+\cos k_{y}\right)\tau^{z}+M\tau^{x}\right]_{\mu\nu}\hat{c}^{\dagger}_{\bm{k}\mu s}\hat{c}_{\bm{k}\nu s}$ (30)
See Fig. 2 for an illustration. In the limit $M\gg|t|$, this electronic structure has two topologically trivial, spin-degenerate, flat bands with width $W\approx 8t^{2}/M$ separated by a band gap $\Delta\approx 2M$. Due to time reversal and spin conservation, as discussed in Sec. V.1, QGN for uniform pairing (p-p channel at $\bm{Q}=\bm{0}$ with $N_{\mu s,\nu s^{\prime}}=\tau^{0}_{\mu\nu}\sigma^{y}_{ss^{\prime}}$) is automatically satisfied. Moreover, since the wavefunctions also obey $\tau^{x}U(\bm{k})=U(\bm{k}+\bm{Q})=U(\bm{Q}-\bm{k})$ with $\bm{Q}=(\pi,\pi)$, we find that QGN is also satisfied in both p-p and p-h channels at this momentum. This can be shown by computing the lowest two eigenvalues of $\Pi^{\bm{Q}}$ in both p-p and p-h channels (Eqs. 5&6), $\varpi^{\bm{Q}}_{0,1}=2\pm(\cos Q_{x}+\cos Q_{y})$ to leading order in $t/M$, so that one branch reaches zero at $\bm{Q}=(0,0)$ and the other at $\bm{Q}=(\pi,\pi)$. Solving the corresponding nesting matrices at $\bm{Q}=(\pi,\pi)$, we find $N_{\mu s,\nu s^{\prime}}=\tau^{x}_{\mu\nu}\sigma^{y}_{ss^{\prime}}$ in the p-p channel, and $N^{(\mathsf{i}=0,x,y,z)}_{\mu s,\nu s^{\prime}}=\tau^{x}_{\mu\nu}\sigma^{(\mathsf{i})}_{ss^{\prime}}$ in the p-h case, where $\sigma^{(\mathsf{i}=0,x,y,z)}$ are Pauli matrices acting on spin indices. QGN thus suggests that, besides the uniform pairing order, the system is simultaneously prone to a singlet PDW, a charge DW, and a spin DW, all at $\bm{Q}=(\pi,\pi)$. [In Sec. V D of the SM, we generalize this construction to show that any single-particle Hamiltonian obeying $Nh(\mathbf{k})N^{\dagger}=h(\mathbf{k}+\mathbf{Q})$ with a unitary nesting matrix $N$ has perfect QGN in the p-h channel at $\bm{Q}$. (Analogous statements hold in the p-p channel.) For this class of systems satisfying QGN, we further show that the Fubini-Study quantum metric is also ‘nested’, $g(\mathbf{k})=g(\mathbf{k}+\mathbf{Q})$, a property that may guide the search of electronic structures with QGN.]
Figure 2: (a) An illustration of the two-orbital model in Eqs. 30 and the band structure. (b) Cooper pair excitation spectrum above the PDW groundstate for the onsite anti-ferromagnetic and inter-orbital attractive interactions in Eq. 31, choosing $V=t=1,M=5$. The analytical results for the artificial setup discussed in the text ($W=0,\Delta=\infty$ with quantum geometry fixed) are shown in blue (for the two bound states) and red (degenerate particle-particle continuum), whereas the numerical results that are accurate to the first order in $W/V$ are plotted in gray. The lowest, condensing Cooper pair band exhibits a zero mode at $\mathbf{P}=(\pi,\pi)$, agreeing with the fact that the PDW OP $\hat{O}^{\text{p-p},\bm{Q}=(\pi,\pi)}$ commutes with the projected Hamiltonian. It can be seen that this band is insensitive to the correction in $W/V$, while the continuum is significantly broadened, swallowing the upper bound state (Leggett mode) band. The condensing Cooper pair mass can be extracted from a quadratic fitting near $(\pi,\pi)$, which we plot with a dashed line; the value of the mass agrees with the result in the flat band limit Eq. 33.
For different suggested orders, we can construct different ideal interactions adopting our systematic scheme in Sec. II. To give an example, below we focus on the PDW case with $\mathbf{Q}=(\pi,\pi)$ where the nesting matrix is $\tau^{x}\sigma^{y}$ corresponding to inter-orbital singlet pairing. We pick a judicious choice of interaction,
$\hat{H}_{\text{int}}=V\sum_{\bm{R}}(-3\hat{n}_{\bm{R},1}\hat{n}_{\bm{R},2}+4\hat{\bm{S}}_{\bm{R},1}\cdot\hat{\bm{S}}_{\bm{R},2}),$ (31)
that uniquely selects the PDW state as the ground state (see Sec. V C of the SM). $\hat{\bm{S}}_{\bm{R},\mu}=\frac{1}{2}\sum_{s}\hat{c}^{\dagger}_{\bm{R},\mu,s}\boldsymbol{\sigma}_{ss^{\prime}}\hat{c}_{\bm{R},\mu,s^{\prime}}$ is the spin operator on layer-$\mu$ and site-$\bm{R}$. The exact ground states are generated by the p-p order parameter (Cooper pair creation operator)
$\hat{O}^{\text{p-p},\dagger}=\sum_{\bm{k}\mu\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu,\uparrow}\tau_{\mu\nu}^{x}\hat{\bar{c}}^{\dagger}_{-\bm{k},\nu,\downarrow}$ (32)
with $\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\nu}=\sum_{\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu}P_{\mu\nu}(\bm{k})$ the projection of the orbital operators onto the flat bands (see Eq. (9)). Projecting $\hat{H}_{\text{int}}$ in Eq. (31) to the flat band Hilbert space results in trivial one-body (Hartree-Fock-type) terms that act as a chemical potential on the flat bands. The triviality of the one-body terms generalizes the “uniform pairing condition” Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) utilized in attractive Hubbard models (see Sec. II D of the SM).
Adopting the methods in Sec. IV, we can exactly solve the few-body excitations in an artificial setup by setting the $W\to 0$ and $\Delta\to\infty$ with the quantum geometry $U_{\mu n}(\bm{k})$ remaining fixed. First, we consider the charge ±$\pm 1$ fermionic quasi-particles above the ground states. We find that these excitations are spin-degenerate, gapped, and flat, with energy $E^{(\pm 1)}(\mathbf{p})=3V$. Next, we consider a Cooper pair excitation. This spectrum is also analytically solvable since the two-particle scattering matrix can be diagonalized by generalizing the approach of Ref. Herzog-Arbeitman et al. (2022a) to extract a low-rank Hamiltonian for the pair binding energies. We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$. The Goldstone mode has dispersion
$E^{(2)}_{0}(\mathbf{P})$ $=6Vg(2+\cos P_{x}+\cos P_{y})=6Vg\frac{\delta\bm{P}^{2}}{2}+\dots$ (33)
to leading order in $t/M$, where the coefficient $g$ is the integrated Fubini-Study metric
$g$ $=\int\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{2}\text{Tr }(\partial_{i}P)^{2}=\frac{t^{2}}{M^{2}}+\dots$ (34)
In the second equality of Eq. 33, we expanded in $\delta\bm{P}=\mathbf{P}-(\pi,\pi)$ to find the inverse Cooper pair mass $\approx\frac{6Vt^{2}}{M^{2}}$. We thus find the quantum geometric origin of PDW in this model is similar to the known case of uniform pairing Peotta and Törmä (2015); however, we note that for general cases, the Cooper pair mass may not directly relate to the value of the metric $g$. The dispersion of the higher bound state (Leggett mode) is identical to Eq. (33) up to a constant shift.
One may be worried about the effects of the actual small but non-zero bandwidth $W\sim t^{2}/M$ in our model. To address this issue, we first note that the quasi-particle excitations are relatively unaffected due to their large gap $\sim V\gg W$. For the two-particle excitations, we check the robustness of the key results by numerically calculating the excitation spectrum (shown in Fig. 2), which is obtained by diagonalizing the perturbed scattering matrix in the presence of a small bandwidth $W$ (compared to the interaction strength $V$) and is accurate to the leading order of $W/V$. We indeed see that although the particle-particle continuum around $6V$ is split by the finite bandwidth and the upper bound state branch merges into the continuum, the dispersion of the lowest branch is relatively unchanged. Especially, it retains its minimum at $(\pi,\pi)$ and the Cooper pair mass does not change significantly from its value in the idealized flat band setup.
After verifying the robustness of the two-particle excitation results in Eq. 33, we finally turn to the estimate of the superfluid stiffness and thus the BKT transition temperature $T_{c}$ of the system in the presence of symmetry lowering terms (see Sec. IV.2). Since for generic fillings, the ODLRO density of the system, $\phi_{0}$, is $\mathcal{O}(1)$, $T_{c}$ is expected to scale as
$T_{c}\sim V\frac{t^{2}}{M^{2}}$ (35)
which applies as long as $\Delta\sim M\gg V\gg W\sim t^{2}/M$. In the further limit $V\gg t$, we see that $T_{c}\gg W^{2}/V$, the conventional strong coupling scaling for single-band systems. This observation further corroborates that the phase stiffness originates from the quantum geometry of the band instead of the (single-particle) kinetic energies. This example (along with similar recent efforts in Refs. Jiang and Barlas (2023); Chen and Huang (2023)) is based on the geometric structure of the flat bands and is distinguished by its analytically controlled, microscopic approach. QGN thus points to a fresh direction in the pursuit of PDW phases, intrinsically distinct from the existing literature Han and Kivelson (2022); Wu et al. (2023a); Liu and Han (2023); Wu et al. (2023b); Jin et al. (2022); Han et al. (2020); Setty et al. (2023); Haldane et al. (2023).
VI Possible Extensions
We note that the form of solvable models presented in Sec. II is not the only possibility for systems with perfect QGN. For instance, consider the one-dimensional SSH chain
$\hat{H}_{0}=\sum_{k\alpha\beta s}t[\tau^{0}-\tau^{z}\cos k+\tau^{y}\sin k]_{\alpha\beta}\hat{c}^{\dagger}_{k,\alpha,s}\hat{c}_{k,\beta,s}$ (36)
which realizes an obstructed atomic limit Song et al. (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al. (2022b); Espinosa-Champo and Naumis (2023). As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al. (2022a). However, we find that this result can be promoted to a singlet $p$-wave OP
$\hat{O}^{\text{p-p},\dagger}=\sum_{k}\sin k\,\gamma^{\dagger}_{k,1,\uparrow}\gamma^{\dagger}_{-k,1,\downarrow}$ (37)
by choosing the interaction taking the form $H_{\text{int}}=U\sum_{R}\hat{S}^{\dagger}_{\bm{R}}\hat{S}_{\bm{R}}\geq 0$ where $\hat{S}_{\bm{R}}$ again satisfies $[\hat{P}\hat{S}_{\bm{R}}\hat{P},\hat{O}^{\text{p-p},\dagger}]=0$ but is non-hermitian. This generalizes the form assumed in Eq. 12 (see Sec. VI of SM for the explicit construction which guarantees trivial one-body terms). Remarkably, zero-energy ground states can still be exactly constructed according to $(\hat{O}^{\text{p-p},\dagger})^{n}\ket{0}$ and proven to be unique using a corollary of Lieb’s well-known theorem Lieb (1989). Although the excitations in this construction are not exactly solvable, we can prove (see Sec. VI of SM) that the charge ±$\pm 1$ excitation spectrum is gapless, as expected for a $p$-wave ground state. This example demonstrates that the possible construction schemes of “ideal” interacting Hamiltonians for systems with QGN have not been exhausted, which we leave for future explorations.
Moreover, it remains to be investigated whether QGN is the necessary condition for a set of flat bands to admit solvable strongly coupled interacting Hamiltonians. Our work advances the exploration Salerno et al. (2023); Huang et al. (2022); Yu et al. (2023); Kashihara et al. (2023); Herzog-Arbeitman et al. (2022); Soldini et al. (2023); Jiang et al. (2023); Yu et al. (2023); Peralta Gavensky et al. (2023); Setty et al. (2023, 2023) of many-body phases enabled by topology and quantum geometry making use of the strongly coupled flat band limit, but it also important to explore how QGN can be combined with Fermi surface nesting for dispersive systems featuring significant quantum geometry.
Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-da Song, and Päivi Törmä for helpful discussions. ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616. BAB was primarily supported by the the Simons Investigator Grant No. 404513. JHA was supported by a Hertz Fellowship. BAB and JHA received additional support from the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, Office of Naval Research (ONR Grant No. N00014-20-1-2303), BSF Israel US foundation No. 2018226 and NSF-MERSEC DMR-2011750, Princeton Global Scholar and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No 101017733 and from the European Research Council (ERC). SAK and ZH were supported, in part, by the Department of Energy, Office of Basic Energy Sciences, under contract DE-AC0276SF00515 at Stanford.

### v2

@@
 The properties of a correlated electronic system are largely determined by its electronic structure.
-One such familiar manifestation is the role of Fermi surface nesting conditions: if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e.
+One such familiar manifestation is the role of Fermi surface nesting conditions Lomer (1962); Fedders and Martin (1966): if one portion of a Fermi surface overlaps with another under a shift by a wavevector $\bm{Q}$ (i.e.
 $\bm{k}\to\bm{k}+\bm{Q}$), the Fermi surface is said to be “perfectly nested” and consequently the system is prone to density wave (DW) order at $\bm{Q}$ with particle-hole (p-h) fermion bilinear order.
@@
 (2022) for typical applications).
-However, their validity is limited to weakly coupled systems, and they completely neglect another important piece of information in the electronic structure - the quantum geometry Resta (2011) and topology Bradlyn et al.
+However, their validity is limited to weakly coupled systems in which the interaction strength is small compared to the bandwidth.
+In general, all forms of non-interacting susceptibilities grow as decreasing $T$ with the same asymptotic behavior $\chi\sim 1/T$ for $T$ greater than the bandwidth, which thus indicates comparable ordering tendencies for all possible order parameters.
+Significantly, this concept, defined solely by the electronic dispersion, overlooks another important piece of information in the electronic structure - the quantum geometry Provost and Vallee (1980); Resta (2011) and topology Bradlyn et al.
 (2017); Po et al.
@@
 (2022); Herzog-Arbeitman et al.
-(2022a); Peotta (2022); Khalaf et al.
+(2022, 2022); Peotta (2022); Khalaf et al.
 (2022); Chan et al.
@@
 (2023); Wu et al.
-(2023); Chen and Huang (2023).
+(2023); Chen and Huang (2023); Salerno et al.
+(2023); Huang et al.
+(2022); Yu et al.
+(2023); Kashihara et al.
+(2023); Soldini et al.
+(2023); Jiang et al.
+(2023); Yu et al.
+(2023); Peralta Gavensky et al.
+(2023); Setty et al.
+(2023); Verma et al.
+(2023); Setty et al.
+(2023a).
 In this paper, we propose such a concept, which we call “Quantum Geometric Nesting” (QGN), that focuses on the geometric aspects of the electronic structure in flat-band systems.
@@
 We will show that these questions can be addressed by examining whether a set of QGN conditions is satisfied.
+Intuitively, these conditions refer to certain forms of ‘overlap’ of the wavefunctions in the flat band subspace at wavevectors related by a shift (p-h case) or inversion (p-p case) in momentum space, akin to the familiar case of Fermi surface nesting.
 For “perfectly nested quantum geometry,” we construct ideal Hamiltonians containing arbitrary combinations of an infinite class of engineered short-range two-body interactions that have solvable ground states with fermion bilinear order of a character determined by the quantum geometry.
@@
 For the p-p case, we further show that these ideal Hamiltonians host a (fine-tuned) pseudospin $SU(2)$ symmetry, whose stiffness can be inferred from the excitation spectra.
-These solvable interacting Hamiltonians can either directly describe realistic systems or serve as starting points for various perturbative studies of additional terms.
-This type of intrinsically strong coupling viewpoint has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al.
+We prove that certain symmetries (e.g.
+chiral or time-reversal) can ensure particular forms of QGN in flat bands.
+Non-symmetry-related cases also exist, such as an engineered model for pair-density-wave order that we construct.
+Extending the construction scheme, we further obtain an SSH-type model with solvable $p$-wave pairing groundstates.
+Indeed, except for symmetry-enforced cases, it is unlikely for a physical system to have exact QGN - in common with the case of Fermi surface nesting.
+However, the use of QGN can be much more versatile: for systems with almost perfect QGN (quantification of the closeness to the ideal nesting will be defined below as ‘nestability’), one may engineer an idealized electronic structure with perfect QGN and then construct ideal interacting Hamiltonians that are close to the physical one.
+The solvability of these ideal models thus makes them excellent starting points for including perturbative corrections to reach the physical Hamiltonian.
+This type of intrinsically strong coupling viewpoint (in the sense that the dominant part of the interactions has been accounted for in the solvable models) has proven fruitful in recent studies of twisted bilayer graphene and other flat-band materials Tovmasyan et al.
 (2016); Kwan et al.
@@
 (2022); Herzog-Arbeitman et al.
-(2022a); Zhang et al.
+(2022); Zhang et al.
 (2022); Herzog-Arbeitman et al.
@@
 (2023); Kwan et al.
-(2023), and it is remarkable that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems.
-More possibilities also exist within the framework of QGN, which are exemplified with an engineered model for pair-density-wave.
-Extending the construction scheme, we further obtain an SSH-type model with solvable $p$-wave pairing groundstates.
+(2023).
+It is noteworthy that perfect QGN is an underlying feature of most of the known examples of solvable models for flat-band systems featuring fermion bi-linear order.
+We hence expect that QGN can greatly facilitate studies within this paradigm, in a way similar to the roles played by ‘ideal band conditions’ Roy (2014); Claassen et al.
+(2015); Mera and Ozawa (2021); Ledwith et al.
+(2022); Wang et al.
+(2021); Wang and Liu (2022); Wang et al.
+(2023); Dong et al.
+(2023) and “vortexability” Ledwith et al.
+(2023) in studies of fractional Chern insulator states.
 We will present the conclusions in the main text and defer the detailed proofs and technical discussions to the corresponding sections in the Supplemental Materials (SM).
@@
 IV).
+A practical recipe for using QGN in perturbative analyses of physical systems is proposed in Sec.
+II.
 Some examples of electronic structures with perfect QGN can be found in Sec.
@@
 I.1 Preliminaries
-We consider an electronic structure encoded in a tight-binding model with charge conservation and translation symmetry among $\mathsf{V}$ unit cells in arbitrary dimension:
+We consider an electronic structure encoded in a tight-binding model  [^1] One could also formulate the model in continuum and work in momentum space throughout the discussion; no conclusions will be affected so long as lattice is not considered.
+Here we start with an explicit tight-binding model to avoid the subtleties when discussing the locality of the interactions and to connect to some existing literature.
+with charge conservation and translation symmetry among $\mathsf{V}$ unit cells in arbitrary dimension:
 $\hat{H}_{0}=\sum_{\bm{R}\bm{R}^{\prime},\mu\nu}t_{\mu\nu}(\bm{R}-\bm{R}^{\prime})\ \hat{c}^{\dagger}_{\bm{R},\mu}\hat{c}_{\bm{R}^{\prime},\nu}$ (1)
 where $\bm{R}$ indicates the unit cell position, Greek letters in lowercase $\mu,\nu,\dots$ denote a “flavor” index which, for simplicity, includes both orbital and spin indices unless otherwise specified.
-The annihilation operators for Bloch states with a given flavor index are expressed as $\hat{c}_{\bm{k},\mu}\equiv\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{R}}\hat{c}_{\bm{R},\mu}/\sqrt{\mathsf{V}}$ (periodic gauge) while the single particle Hamiltonian is diagonalized in a band basis,
+The annihilation operators for Bloch states with a given flavor index are expressed as $\hat{c}_{\bm{k},\mu}\equiv\sum_{\bm{R}}\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{R}}\hat{c}_{\bm{R},\mu}/\sqrt{\mathsf{V}}$ in the periodic embedding.
+[^2] QGN can be defined for any embedding of the intra-unit-cell positions $\{\bm{r}_{\mu}\}$, but the periodic embedding $\{\bm{r}_{\mu}=\bm{0}\}$ is a special one since if a certain QGN condition is satisfied for any embedding, it is also satisfied in the periodic embedding.
+It thus suffices to always work with periodic embedding.
+In the sense that the embedding can be fixed, QGN is an embedding-independent concept, which is necessary for QGN to be a physical concept since embedding is purely a redundancy in the definition of Bloch basis Huhtinen et al.
+(2022).
+In momentum space, the single particle Hamiltonian is diagonalized with a band basis
 $\hat{H}_{0}$ $=\sum_{\bm{k},n}\epsilon_{n}(\bm{k})\hat{\gamma}^{\dagger}_{\bm{k},n}\hat{\gamma}_{\bm{k},n},\quad\hat{\gamma}_{\bm{k},n}=U_{n\mu}^{\dagger}(\bm{k})\hat{c}_{\bm{k},\mu},$ (2)
@@
 We will quantify the degree of QGN for flat-band systems by utilizing the geometric information in $U_{\mu n}(\bm{k})$.
-To do so, we consider an idealized setup where there are $\mathsf{N}_{\text{flat}}$ nearly degenerate and flat bands isolated from all the other bands by a large gap $\Delta$, such that the relevant interaction strength $V$ is small compared to $\Delta$ but large compared to the flat bands’ bare bandwidth $W$,
+To do so, we consider an idealized setup (illustrated in Fig.
+1) where there are $\mathsf{N}_{\text{flat}}$ nearly degenerate and flat bands isolated from all the other bands by a large gap $\Delta$, such that the relevant interaction strength $V$ is small compared to $\Delta$ but large compared to the flat bands’ bare bandwidth $W$,
 $W\ll V\ll\Delta.$ (3)
@@
 I.2 General Definitions
-In terms of the gauge-invariant matrices $P$ and $Q$, the key information of QGN at a given wavevector $\bm{Q}$ for p-p or p-h channel is encoded in the linear operator
+In terms of the gauge-invariant projection matrices $P$ and $Q$, the key information of QGN at a given wavevector $\bm{Q}$ for p-p or p-h channel is encoded in the linear operator
 $\Pi^{\text{p-p},\bm{Q}}_{\mu^{\prime}\nu^{\prime};\mu\nu}=\frac{1}{\mathsf{V}}\sum_{\bm{k}}\left[P^{\star}_{\mu^{\prime}\mu}(\frac{\bm{Q}}{2}+\bm{k})Q_{\nu\nu^{\prime}}(\frac{\bm{Q}}{2}-\bm{k})+(P\leftrightarrow Q)\right],$ (5)
 $\Pi^{\text{p-h},\bm{Q}}_{\mu^{\prime}\nu^{\prime};\mu\nu}=\frac{1}{\mathsf{V}}\sum_{\bm{k}}\left[P_{\mu^{\prime}\mu}(\bm{k}+\frac{\bm{Q}}{2})Q_{\nu\nu^{\prime}}(\bm{k}-\frac{\bm{Q}}{2})+(P\leftrightarrow Q)\right].$ (6)
-Viewed as a matrix (with $\mu\nu$ a grouped index and $\mu^{\prime}\nu^{\prime}$ the other), $\Pi$ is hermitian and positive semi-definite and can be diagonalized with non-negative eigenvalues.
+where means complex conjugation.
+When viewed as a matrix (with $\mu\nu$ a grouped index and $\mu^{\prime}\nu^{\prime}$ the other), $\Pi$ is hermitian and positive semi-definite and can be diagonalized with non-negative eigenvalues.
 By definition, an electronic structure has “perfect QGN” at a wavevector $\bm{Q}$ in the p-p or p-h channel when the corresponding $\Pi$ has eigenvalue(s) equal to 00.
-In other words, a perfect QGN is satisfied if there exist a $\mathsf{N}\times\mathsf{N}$ matrix $N^{\bm{Q}}_{\mu\nu}$, which we call the “nesting matrix”, satisfying (for any $\mu^{\prime}\nu^{\prime}$)
+In other words, a perfect QGN is satisfied if there exists a $\mathsf{N}\times\mathsf{N}$ matrix $N^{\bm{Q}}_{\mu\nu}$, which we call the “nesting matrix”, satisfying (for any $\mu^{\prime}\nu^{\prime}$)
 $\sum_{\mu\nu}\Pi_{\mu^{\prime}\nu^{\prime};\mu\nu}^{\bm{Q}}N^{\bm{Q}}_{\mu\nu}=0.$ (7)
@@
 $\hat{O}^{\text{p-h},\bm{Q}}$ $\equiv\frac{1}{\mathsf{V}}\sum_{\bm{k};n,m\leq\mathsf{N}_{\text{flat}}}F^{\text{p-h},\bm{Q}}_{nm}(\bm{k})\hat{\gamma}^{\dagger}_{\bm{k}+\bm{Q}/2,n}\hat{\gamma}_{\bm{k}-\bm{Q}/2,m}$ (9)
-where $F$ is the form factor
+where $F(\bm{k})$ is the form factor in the band basis corresponding to $N^{\bm{Q}}$ being the unprojected form factor in the original orbital basis,
 $F^{\text{p-p},\bm{Q}}_{nm}(\bm{k})\equiv\sum_{\mu\nu}U_{\mu n}(\frac{\bm{Q}}{2}+\bm{k})N^{\text{p-p},\bm{Q}}_{\mu\nu}U_{\nu m}(\frac{\bm{Q}}{2}-\bm{k})$ (10)
 $F^{\text{p-h},\bm{Q}}_{nm}(\bm{k})\equiv\sum_{\mu\nu}U^{\dagger}_{n\mu}(\bm{k}+\frac{\bm{Q}}{2})N^{\text{p-h},\bm{Q}}_{\mu\nu}U_{\nu m}(\bm{k}-\frac{\bm{Q}}{2}).$ (11)
-One important property (which can serve as an equivalent definition) of QGN is that for arbitrary $\bm{k}$, the form factor $F_{nm}(\bm{k})$ defined by a nesting matrix is a block-diagonal matrix that disconnects flat band indices from the others, i.e.
+One important property (which can actually serve as an equivalent definition) of QGN is that for arbitrary $\bm{k}$, the form factor $F_{nm}(\bm{k})$ defined by a nesting matrix is a block-diagonal matrix that disconnects flat band indices from the others, i.e.
 $F_{nm}(\bm{k})=0$ as long as $n,m$ are not simultaneously $\leq\mathsf{N}_{\text{flat}}$ or $>\mathsf{N}_{\text{flat}}$.
+In other words, QGN implies that the wavefunctions of the flat bands at two momenta related by a shift (p-h case) or inversion (p-p case) are fully compatible, in the sense $N^{\bm{Q}}$ maps the single-particle flat band null spaces to each other.
+As a direct result, the static susceptibility of the OP defined in Eqs.
+8&9 can be shown (see Sec.
+I C of SM) to strictly saturate an upper bound when the corresponding QGN is satisfied, suggesting the harmony between the quantum geometry and the OP in the ideal circumstances.
+We will further show in Sec.
+II that an infinite class of interacting solvable models can be constructed, which host ground states with the corresponding order.
 For QGN to be non-trivial in the p-p channel, it is additionally required that $N^{\bm{Q}}$ cannot be symmetric since $\hat{O}^{\text{p-p},\bm{Q}}$ vanishes in this case.
 (It is proven by explicit construction that valid, anti-symmetric solutions can indeed exist for $\Pi$.) On the other hand, the identity matrix is always an allowed solution for Eq.
-7 in the p-h case for $\bm{Q}=\bm{0}$, but this is a trivial one in the sense that the resulting $\hat{O}^{\text{p-h},\bm{Q}=\bm{0}}$ is simply the density operator not associated with any broken symmetry.
-Excluding all the unwanted solutions, for a generic flat-band system, the value of the smallest eigenvalue of $\Pi^{\bm{Q}}$, $\varpi_{0}^{\bm{Q}}$, is a natural dimensionless measure of “nestability” that quantifies the closeness of a system to perfect QGN in the corresponding channel at $\bm{Q}$  [^2] We note that these concepts are spiritually similar to the “ideal band condition” Roy (2014); Claassen et al.
+7 in the p-h case for $\bm{Q}=\bm{0}$, but this is a trivial one in the sense that the resulting $\hat{O}^{\text{p-h},\bm{Q}=\bm{0}}$ is simply the total density operator not associated with any broken symmetry.
+Excluding all the unwanted solutions, for a generic flat-band system, the value of the minimal eigenvalue of $\Pi^{\bm{Q}}$, $\varpi_{0}^{\bm{Q}}$, is a natural dimensionless measure of “nestability” that quantifies the ‘distance’ of a system to perfect QGN in the corresponding channel at $\bm{Q}$  [^4] We note that these concepts are spiritually similar to the “ideal band condition” Roy (2014); Claassen et al.
 (2015); Mera and Ozawa (2021); Ledwith et al.
@@
 triplet SC in spin invariant systems.
+We mention that each satisfied QGN condition allows for constructing a generalized quantum geometry in the sense that it defines a notion of distance in momentum space and thus a generalized metric tensor, which can be related to the second spatial moment of the OP density bare correlation functions originating from the flat bands.
+The familiar definition of quantum distance $d(\bm{k},\bm{k}^{\prime})\equiv\sqrt{\mathsf{N}_{\text{flat}}-{\rm tr}{\left[P(\bm{k})P(\bm{k}^{\prime})\right]}}$ Provost and Vallee (1980); Resta (2011) using the Hilbert-Schmidt norm turns out to be one special case in this class of constructions, which is always valid since it corresponds to the ‘trivial QGN’ we mentioned above with $N^{\text{p-h},\bm{Q}=\bm{0}}=\mathbbm{1}$.
+We refer interested readers to Sec.
+I D of SM for formal definitions and discussions.
 II Ideal interacting Hamiltonians
@@
 14 exists.
-However, when $N$ and $U$ in Eq.
-8/9 derive from a band geometry with perfect QGN, there is an infinite class of solutions that take a special flavor-momentum separable form
-$\hat{S}^{(I)}_{\bm{R}}=\sum_{\bm{R}_{1}\bm{R}_{2};\mu\nu}s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}B^{(I)}_{\mu\nu}\hat{c}^{\dagger}_{\bm{R}+\bm{R}_{1},\mu}\hat{c}_{\bm{R}+\bm{R}_{2},\nu},$ (15)
-where $B^{(I)}$ is a hermitian matrix and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=[s^{(I)}_{\bm{R}_{2},\bm{R}_{1}}]^{\star}$ is the spatial coeffient.
+However, when the OP derives from a band geometry with perfect QGN, there is (at least) an infinite class of solutions that take a special flavor-space separable form
+$\hat{S}^{(I)}_{\bm{R}}=$ $\sum_{\bm{R}_{1}\bm{R}_{2};\mu\nu}A^{(I)}(\bm{R}_{1},\bm{R}_{2})B^{(I)}_{\mu\nu}\hat{c}^{\dagger}_{\bm{R}+\bm{R}_{1},\mu}\hat{c}_{\bm{R}+\bm{R}_{2},\nu},$ (15)
+where $A^{(I)}(\bm{R}_{1},\bm{R}_{2})=[A^{(I)}(\bm{R}_{2},\bm{R}_{1})]^{\star}$ is a spatial coefficient and $B^{(I)}$ is a hermitian matrix acting on orbital indices.
 In order for Eq.
-14 to hold, the constraints on the coefficient $s^{(I)}$ and matrix $B^{(I)}$ are rather loose, in that they only need to obey
-$s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=\left[s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}\right]^{\star}\mathrm{e}^{-\mathrm{i}(\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}},$ (16)
+14 to hold, the constraints on the coefficient $A^{(I)}$ and matrix $B^{(I)}$ are not overly restrictive, in that they only need to obey
+$A^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=\left|A^{(I)}_{\bm{R}_{1},\bm{R}_{2}}\right|\mathrm{e}^{-\mathrm{i}(\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}/2},$ (16)
 $N^{\text{p-p},\bm{Q}}\cdot B^{(I)}+[B^{(I)}]^{T}\cdot N^{\text{p-p},\bm{Q}}=0,$ (17)
 for the p-p case, and
-$s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}=0\ \ \text{if}\ \ (\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}\neq 0\mod 2\pi,$ (18)
+$A^{(I)}(\bm{R}_{1},\bm{R}_{2})=0\ \ \text{if}\ \ (\bm{R}_{1}-\bm{R}_{2})\cdot\bm{Q}\neq 0\text{ mod }2\pi,$ (18)
 $N^{\text{p-h},\bm{Q}}\cdot B^{(I)}-B^{(I)}\cdot N^{\text{p-h},\bm{Q}}=0.$ (19)
@@
 V.) In particular, the identity matrix is always a solution for $B$ in the p-h case, which makes $\hat{S}_{\bm{R}}$ a local density operator for the unit cell at $\bm{R}$.
-Clearly, even given all the constraints, we are still left with an infinite class of possible interaction terms with considerable leeway, at least in choosing $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ and $s^{(I)}_{\bm{R}_{1},\bm{R}_{2}}$, which thus could substantially overlap with realistic interactions.
+Clearly, even given all the constraints, we are still left with an infinite class of possible interaction terms with considerable leeway, at least in choosing $V_{IJ}(\bm{R}-\bm{R}^{\prime})$ and $A^{(I)}(\bm{R}_{1},\bm{R}_{2})$.
+Although the above construction scheme could encompass many realistic forms of interactions, it is admittedly unlikely that the electronic structure and interactions of any physical system are perfectly ideal.
+Therefore, the merit of QGN should be viewed in a more practical way: it can help identify solvable limits of physical flat band systems of interest.
+Specifically, we are proposing the following recipe for the use of QGN in the study of a physical Hamiltonian $\hat{H}_{\text{phys}}$: First, one computes its electronic structure and examine whether there is a channel and a wavevector on which QGN is nearly satisfied (i.e.
+the corresponding $\varpi_{0}$ is small compared to $1$).
+If there is indeed such a nearly nested possibility, one can consider its perfect nesting limit by slightly modifying the electronic structure either numerically or analytically (through recognizing certain approximate symmetries, e.g.
+those discussed below in Sec.
+V).
+Then, one can further search in the space of all possible ideal solvable interacting Hamiltonians validated by the QGN, to find one that is closest to the physical Hamiltonian, $\hat{H}_{\text{ideal}}$.
+This can be systematically achieved by minimizing the norm of the difference between them, $||\delta\hat{H}||=||\hat{H}_{\text{ideal}}-\hat{H}_{\text{phys}}||$.
+A small value of this operator distance justifies perturbative analyses starting from the ideal solvable limit.
+The flexibility of our construction allows the dominant part of many interactions to be captured.
+Furthermore, we remark that these nearby solvable limits can also facilitate numerical studies of $\hat{H}_{\text{phys}}$, in that one can benchmark the validity of the methods being used at these limits, and confirm the phase being observed by checking whether there is an intervening phase transition as interpolating between $\hat{H}_{\text{phys}}$ and $\hat{H}_{\text{ideal}}$.
 One caveat on the terminology should be noted: After projecting onto $\mathcal{H}_{\text{flat}}$, there are terms quadratic in fermion operators, besides the quartic ones that are conventionally called “interactions”.
@@
 (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al.
-(2022a).
+(2022).
 It is possible to generalize this condition, which we will exemplify in Sec.
@@
 II D of SM.
+We remark that the building blocks of the ideal Hamiltonians, $\{\hat{S}^{I}_{\bm{R}}\}$, do not necessarily mutually commute.
+As we will see in the next section, the solvability of these interactions results from the fact that they all commute with the order parameter and share certain eigenstates.
+The ideal interacting Hamiltonians are in this sense ‘frustration-free’.
 III Solvable ground state
@@
 Thus, all $|\Psi_{\text{SC}}(\theta)\rangle$ with different $\theta$ are simultaneous eigenstate of $\hat{S}^{(I)}_{\bm{R}}$ with the same eigenvalue.
-The corresponding ground state of $\hat{H}_{\text{int}}$ with a fixed (even) particle number $\mathsf{N}_{e}$ can be constructed as
+The corresponding ground state of $\hat{H}_{\text{int}}$ with an arbitrary fixed (even) particle number $\mathsf{N}_{e}$ can be constructed as
 $|\Psi_{\text{SC}}(\mathsf{N}_{e})\rangle\equiv\int_{0}^{2\pi}\mathrm{e}^{-\mathrm{i}\theta\mathsf{N}_{e}/2}|\Psi_{\text{SC}}(\theta)\rangle\ .$ (22)
 In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector.
+This implies that the system is solvable only at certain rational filling fractions of the flat bands with specific values determined by the OP.
 It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure.
@@
 (2021); Herzog-Arbeitman et al.
-(2022a); Vafek and Kang (2021) and thus the composition does not changed after the commutation.
+(2022); Vafek and Kang (2021) and thus the composition does not changed after the commutation.
 For example, for two-particle (Cooper pair) excitations,
@@
 (2021); Herzog-Arbeitman et al.
-(2022a).
+(2022).
 For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian.
@@
 attractive Hubbard Herzog-Arbeitman et al.
-(2022a)), the quasi-particle spectrum can be even analytically solvable.
+(2022)), the quasi-particle spectrum can be even analytically solvable.
 IV.2 Pseudo-spin $SU(2)$ symmetry of the p-p case and the phase stiffness
@@
 (2016); Herzog-Arbeitman et al.
-(2022a) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al.
+(2022) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al.
 (2023); Liang et al.
-(2017); Verma et al.
+(2017a); Verma et al.
 (2021); Hazra et al.
@@
 (2020); Herzog-Arbeitman et al.
-(2022b); Huhtinen et al.
+(2022); Huhtinen et al.
 (2022); Herzog-Arbeitman et al.
-(2022a).
+(2022).
 V.2 Chiral symmetric flat bands
@@
 (2021).
+In the language of QGN, it is now clear that these proposed solvable interacting Hamiltonians belong to a larger class, which is systematically constructible using the method in Sec.
+II for each possible ordering.
 We also note that there is a large class of electronic structures that feature flat bands protected by chiral symmetry Călugăru et al.
@@
 (2016); Herzog-Arbeitman et al.
-(2022a) utilized in attractive Hubbard models (see Sec.
+(2022) utilized in attractive Hubbard models (see Sec.
 II D of the SM).
@@
 Herzog-Arbeitman et al.
-(2022a) to extract a low-rank Hamiltonian for the pair binding energies.
+(2022) to extract a low-rank Hamiltonian for the pair binding energies.
 We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$.
@@
 This observation further corroborates that the phase stiffness originates from the quantum geometry of the band instead of the (single-particle) kinetic energies.
-This example (along with similar recent efforts in Refs.
-Jiang and Barlas (2023); Chen and Huang (2023)) is based on the geometric structure of the flat bands and is distinguished by its analytically controlled, microscopic approach.
+This example (along with similar efforts in Refs.
+Liang et al.
+(2017b); Jiang and Barlas (2023); Chen and Huang (2023)) is based on the geometric structure of the flat bands and is distinguished by its analytically controlled, microscopic approach.
 QGN thus points to a fresh direction in the pursuit of PDW phases, intrinsically distinct from the existing literature Han and Kivelson (2022); Wu et al.
@@
 (2020); Setty et al.
-(2023); Haldane et al.
+(2023b); Haldane et al.
 (2023).
@@
 (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al.
-(2022b); Espinosa-Champo and Naumis (2023).
+(2022); Espinosa-Champo and Naumis (2023).
 As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al.
-(2022a).
+(2022).
 However, we find that this result can be promoted to a singlet $p$-wave OP
 $\hat{O}^{\text{p-p},\dagger}=\sum_{k}\sin k\,\gamma^{\dagger}_{k,1,\uparrow}\gamma^{\dagger}_{-k,1,\downarrow}$ (37)
-by choosing the interaction taking the form $H_{\text{int}}=U\sum_{R}\hat{S}^{\dagger}_{\bm{R}}\hat{S}_{\bm{R}}\geq 0$ where $\hat{S}_{\bm{R}}$ again satisfies $[\hat{P}\hat{S}_{\bm{R}}\hat{P},\hat{O}^{\text{p-p},\dagger}]=0$ but is non-hermitian.
+by choosing the interaction taking the form $H_{\text{int}}=U\sum_{R}\hat{S}^{\dagger}_{R}\hat{S}_{R}\geq 0$ where $\hat{S}_{R}$ again satisfies $[\hat{P}\hat{S}_{R}\hat{P},\hat{O}^{\text{p-p},\dagger}]=0$ but is non-hermitian.
+Explicitly, we choose $\hat{S}_{R}$ to take the form
+$\hat{S}_{R}$ $=\sum_{\alpha\beta,\sigma}\sum_{\Delta R=-2}^{2}s^{z}_{\sigma}\tau_{\alpha\beta}(\Delta R)\hat{c}^{\dagger}_{R+\Delta R,\alpha,\sigma}\hat{c}_{R,\beta,\sigma}$ (38)
+where $s^{z}_{\uparrow/\downarrow}=\pm 1$ and $\tau(0)=\tau_{2}/4,\tau(\pm 1)=\mp i\tau_{0}/4,\tau(\pm 2)=(-\tau_{2}\pm i\tau_{3})/8$.
 This generalizes the form assumed in Eq.
@@
 Moreover, it remains to be investigated whether QGN is the necessary condition for a set of flat bands to admit solvable strongly coupled interacting Hamiltonians.
-Our work advances the exploration Salerno et al.
-(2023); Huang et al.
-(2022); Yu et al.
-(2023); Kashihara et al.
-(2023); Herzog-Arbeitman et al.
-(2022); Soldini et al.
-(2023); Jiang et al.
-(2023); Yu et al.
-(2023); Peralta Gavensky et al.
-(2023); Setty et al.
-(2023, 2023) of many-body phases enabled by topology and quantum geometry making use of the strongly coupled flat band limit, but it also important to explore how QGN can be combined with Fermi surface nesting for dispersive systems featuring significant quantum geometry.
+It is interesting to explore whether new notions in quantum geometry Bouhon et al.
+(2023) can help discover more general conditions of QGN.
+Our work advances the exploration of many-body phases enabled by topology and quantum geometry making use of the strongly coupled flat band limit, but it also important to explore how QGN can be combined with Fermi surface nesting for dispersive systems featuring significant quantum geometry.
 Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-da Song, and Päivi Törmä for helpful discussions.
+We also thank anonymous reviewers for inspiring questions and comments.
 ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616.

### v3

@@
 (2022); Herzog-Arbeitman et al.
-(2022, 2022); Peotta (2022); Khalaf et al.
+(2022a, 2024a); Peotta (2022); Khalaf et al.
 (2022); Chan et al.
@@
 (2023); Herzog-Arbeitman et al.
-(2023); Jia et al.
+(2024b); Jia et al.
 (2024); Wu et al.
@@
 (2022); Herzog-Arbeitman et al.
-(2022); Zhang et al.
+(2022a); Zhang et al.
 (2022); Herzog-Arbeitman et al.
@@
 (2020).
-In the p-h channel, if $\bm{Q}=\bm{0}$, non-trivial QGN ($N^{\text{p-h},\bm{0}}\neq\mathbbm{1}$) leads to flavor-symmetry-breaking orders, including flavor ferromagnetism and inter-flavor coherent states; the $\bm{Q}\neq\bm{0}$ case promotes those orders to various forms of flavor DWs (e.g.
-spin DWs), with an additional possibility of pure charge DW order for $N^{\text{p-h},\bm{Q}}=\mathbbm{1}$.
+In the p-h channel, if $\bm{Q}=\bm{0}$, non-trivial QGN ($N^{\text{p-h},\bm{0}}\neq\mathbbm{1}$) leads to flavor-symmetry-breaking orders, including flavor ferromagnetism and inter-flavor coherent states; the $\bm{Q}\neq\bm{0}$ cases include various forms of DWs (e.g., spin, charge, or current/flux DWs).
 A system can satisfy multiple QGN conditions in different channels and/or at different $\bm{Q}$s.
@@
 (2016) renders these terms trivial, and it was found that this condition can be guaranteed simply by certain point-group symmetries Herzog-Arbeitman et al.
-(2022).
+(2022a).
 It is possible to generalize this condition, which we will exemplify in Sec.
@@
 This implies that the system is solvable only at certain rational filling fractions of the flat bands with specific values determined by the OP.
-It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure.
+In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
+It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure (corresponding to DW).
 Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice.
-In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
+It is, in principle, possible to have $\varpi^{\bm{Q}}_{0}\approx 0$ in a continuous range of $\bm{Q}$, such that an incommensurate DW order is suggested.
+In such a case, the properties of each incommensurate $\bm{Q}$ should be approached to a good approximation from a nearby commensurate value.
 While the four-fermion term in $H_{\text{int}}$ is translationally invariant, the single particle term is only translationally invariant when all $\langle\hat{S}^{(I)}_{\bm{R}}\rangle$ are.
@@
 In general, however, certain choices of $\hat{H}_{\text{int}}$ may explicitly break translational symmetry, but still preserve other symmetries that are broken spontaneously by $|\Psi\rangle$.
+It should be noted that despite the spiritual similarity to the mean field theories, the Slater determinant states solved by the above approach are the exact many-body ground states of the ideal systems instead of just proxy states capturing the qualitative properties.
+The perturbations away from those ideal limits can be accounted for systematically since not only the ground states but also many aspects of the excitation spectrum can be solved, as we discuss below.
 IV Excitations in the ideal models
@@
 (2021); Herzog-Arbeitman et al.
-(2022); Vafek and Kang (2021) and thus the composition does not changed after the commutation.
+(2022a); Vafek and Kang (2021) and thus the composition does not changed after the commutation.
 For example, for two-particle (Cooper pair) excitations,
@@
 (2021); Herzog-Arbeitman et al.
-(2022).
+(2022a).
 For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian.
@@
 attractive Hubbard Herzog-Arbeitman et al.
-(2022)), the quasi-particle spectrum can be even analytically solvable.
+(2022a)), the quasi-particle spectrum can be even analytically solvable.
 IV.2 Pseudo-spin $SU(2)$ symmetry of the p-p case and the phase stiffness
@@
 Our approach provides a direct estimate of the superfluid stiffness of nearly ideal systems, providing complementary estimates to generic upper bounds Hazra et al.
-(2019); Hofmann et al.
-(2022); Mao and Chowdhury (2023a, b).
+(2019a); Hofmann et al.
+(2022); Mao and Chowdhury (2023, 2024).
 V Examples of electronic structures with perfect QGN
+Here, we present several examples of QGN. More detailed discussions, as well as another constructed model for a current density wave, can be found in Sec. V of the SM.
 V.1 Spinful time-reversal symmetric flat bands
@@
 (2016); Herzog-Arbeitman et al.
-(2022) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al.
+(2022a) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al.
 (2023); Liang et al.
@@
 (2021); Hazra et al.
-(2018); Xie et al.
+(2019b); Xie et al.
 (2020); Herzog-Arbeitman et al.
-(2022); Huhtinen et al.
+(2022b); Huhtinen et al.
 (2022); Herzog-Arbeitman et al.
-(2022).
+(2022a).
 V.2 Chiral symmetric flat bands
@@
 (2016); Herzog-Arbeitman et al.
-(2022) utilized in attractive Hubbard models (see Sec.
+(2022a) utilized in attractive Hubbard models (see Sec.
 II D of the SM).
@@
 Herzog-Arbeitman et al.
-(2022) to extract a low-rank Hamiltonian for the pair binding energies.
+(2022a) to extract a low-rank Hamiltonian for the pair binding energies.
 We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$.
@@
 (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al.
-(2022); Espinosa-Champo and Naumis (2023).
+(2022b); de Jesús Espinosa-Champo and Naumis (2023).
 As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al.
-(2022).
+(2022a).
 However, we find that this result can be promoted to a singlet $p$-wave OP
@@
 Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-Da Song, and Päivi Törmä for helpful discussions.
-We also thank anonymous reviewers for inspiring questions and comments.
 ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616.